What is the dividend yield for Company?

Shares of Common stock 540,000

Shares of Preferred stock 112,000

Common Stock Dividends $2,197,800

Preferred Stock Dividends $1,680,000

Market Price share of common stock $11

0.28

0.37

1.36

1.78
Answer #1

Dividend per share of common stock

= 2,197,800/540,000 = 4.07 per share

Dividend yield = Dividend per share/market price per share

= 4.07/11

= 0.37